# Product of Experts for Visual Generation

### v1

@@ Abstract
Modern neural models capture rich priors and have complementary knowledge over shared data domains, e.g., images and videos. Integrating diverse knowledge from multiple sources—including visual generative models, visual language models, and sources with human-crafted knowledge such as graphics engines and physics simulators—remains under-explored. We propose a Product of Experts (PoE) framework that performs inference-time knowledge composition from heterogeneous models. This training-free approach samples from the product distribution across experts via Annealed Importance Sampling (AIS). Our framework shows practical benefits in image and video synthesis tasks, yielding better controllability than monolithic methods and additionally providing flexible user interfaces for specifying visual generation goals.  [^1] Project page: https://product-of-experts.github.io/.
@@ 1 Introduction
Visual generation requires diverse expertise. State-of-the-art image and video generation models have shown remarkable success in realistic appearance modeling, but still have limitations in producing physically accurate outputs or following complex user instructions. The former would benefit from domain expertise embedded in physics simulators, and the latter from semantic knowledge in large language models (LLMs) and visual language models (VLMs). However, training a single, monolithic model to absorb these varying knowledge sources (massive texts and visual datasets, simulation trajectories, etc.) is often impractical due to misaligned data distributions and training objectives.
We propose to address the challenge by integrating knowledge across diverse models at inference time. Specifically, we formulate visual generation in an inference-time Product of Experts (PoE) framework that aims to synthesize visual content leveraging specialized, pre-existing knowledge sources du2024compositional , including visual generative priors, discriminative rewards from LLMs/VLMs, and rule-based knowledge from physics simulators.
A naïve approach for incorporating multiple constraints is rejection sampling: generate candidate samples (or complete trajectories) using a base expert model, then discard failing ones. However, this is computationally prohibitive in high-dimensional visual synthesis due to vanishing acceptance rates. To improve sample efficiency, we utilize an annealed sampling procedure neal2001annealed with Sequential Monte Carlo (SMC) doucet2001introduction . Instead of inefficient post-hoc filtering, we propose to implement local (spatial for images and spatio-temporal for videos) updates from both generative and discriminative experts over many sampling steps, effectively steering the generation towards valid outputs.
The proposed framework can generate scenes better aligned with user inputs by explicitly integrating complementary expert knowledge sources without costly retraining. Its compositionally provides critical sample efficiency that enables image generation following complex text instructions. It also naturally yields a flexible user interface for defining complex generative goals by selecting and configuring experts, e.g., allowing users to specify the pose and motion trajectory of an object of interest, insert it into an existing image, and animate the full scene.
Overall, our contributions in this work are threefold:
@@
    We formulate visual synthesis tasks under constraints from heterogeneous generative and discriminative models in a unified Product-of-Experts (PoE) framework.
@@
    We propose a practical, sample-efficient framework implementation with Annealed Importance Sampling (AIS) and Sequential Monte Carlo (SMC), with instantiated experts ranging across visual generative models, VLMs, and non-neural physics simulators.
@@ 2 Related Works
Our work is related to recent work on compositional generative modeling, where multiple generative models are combined to jointly generate data samples (du2020compositional, ; garipov2023compositional, ; du2023reduce, ; mahajan2024compositional, ; du2024compositional, ; bradley2025mechanisms, ; thornton2025composition, ; gaudi2025coind, ). Compositional models have been previously applied in visual content generation (liu2022compositional, ; bar2023multidiffusion, ; zhang2023diffcollage, ; yang2024probabilistic, ; su2024compositional, ; li2022composing, ), but such applications typically focus on multiple homogeneous generator experts, or, in the case of li2022composing , a single generator expert with multiple discriminative experts. In contrast, our approach provides a general probabilistic framework through which many heterogeneous generative and discriminative experts can be jointly combined in generation.
@@
Our work is further related to recent work steering generative models with discriminative models and reward functions grathwohl2019your ; dhariwal2021diffusion ; chung2022score ; song2023loss ; bansal2023universal ; he2023manifold ; ye2024tfg ; rout2025semantic . Methods typically use gradient descent on a discriminative reward grathwohl2019your ; dhariwal2021diffusion ; bansal2023universal , which assumes dense, differentiable gradients on the reward and are often approximate samplers that do not support sparse rewards models, e.g., VLMs. A recent branch of work has illustrated how Sequential Monte Carlo can be used for more exact reward steering (wu2023practical, ; zhao2024probabilistic, ). Our work builds on top of this formulation and generalizes it to sample from a composition of generative and discriminative experts.
@@
Our framework integrates both the data-driven knowledge from neural models and the domain-specific knowledge from non-neural models, e.g., physical simulators, allowing for more controllable, physically accurate video generation than end-to-end video generation models brooks2024video ; wan2025 , while alleviating the need for tedious 3D scene setups in traditional graphics rendering pipelines. Recent works have explored such direction typically convert physical simulator outputs into a specific type of conditional signals such as optical flow burgert2025go ; montanaro2024motioncraft ; yang2025towards and point tracking gu2025das , or rely on a single pre-trained video generation model liu2024physgen ; chen2025physgen3d ; tan2024physmotion for video-domain prior, while our framework allows for the integration of various signals available from physical simulation, providing improved flexibility and more complete specifications of controls.
3 The Product of Experts (PoE) Framework
We aim to generate complex scenes $x\in\mathbb{R}^{d}$ by combining the rich priors inside both generative models (such as autoregressive, diffusion models, and control signals converted from classical physics simulators) as well as discriminative models (such as VLMs), with examples detailed in Section 4. We formulate this composition of models in a probabilistic manner, where generative models are represented as probabilistic priors over scenes $\{p^{(i)}(\cdot)\}_{i=1}^{N}$. Discriminative models are represented as constraint functions, $\{r^{(j)}(\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}\}_{j=1}^{M}$, each assigns a scalar score that represents the degree to which a scene $x$ matches the constraint encoded in the model. Each constraint function is converted into an unnormalized probability distributions through the Boltzmann distribution, $q^{(j)}:=e^{r^{(j)}}$.
To combine generative and discriminative models $\{p^{(i)}(\cdot)\}_{i=1}^{N}$ and $\{q^{(j)}(\cdot)\}_{j=1}^{M}$, we sample from the product distribution
@@ 3 The Product of Experts (PoE) Framework
where each generative expert $p^{(i)}(x)$ and discriminative expert $q^{(j)}(x)$ are defined over parts of the scene $x$. The product distribution $p(x)$ has high likelihood at scenes $x$ where each part of the scene is high likelihood under both $p^{(i)}(x)$ and $q^{(j)}(x)$ hinton2002training ; du2020compositional ; du2024compositional , allowing us to compose together the priors in both generative and discriminative models.
However, sampling from Eq. 1 is challenging, as we need to search for a sample $x$ that satisfies constraints across all experts. Below, we will discuss how we can efficiently and effectively sample from the Eq. 1 for high fidelity visual generation. We first discuss how we can sample from multiple generative experts in Section 3.1 and then discuss how to sample jointly from multiple discriminative and generative experts in Section 3.2, followed by a practical implementation in Section 3.3.
@@ 3.1 Sampling from Generative Experts
To sample from the above distribution, we can use Markov chain Monte Carlo (MCMC) robert1999monte , where we can iteratively refine samples based on their likelihood under the product distribution. In discrete settings, a variant of Gibbs sampling may be used, while in continuous domains, Langevin sampling based on the gradient of log density of each distribution may be used welling2011bayesian .
A central issue, however, is that MCMC is a local refinement procedure, where it can take an exponentially long time for the sampling procedure to mix and find a high likelihood scene $x$ from the product distribution robert1999monte . To effectively sample from product of generative experts, we propose to use Annealed Importance Sampling (AIS) neal2001annealed , where we construct of a sequence of $T$ distributions, $\{p_{t}(x)\}_{t=1}^{T}$, where $p_{T}(x)$ corresponds to a smooth, easy-to-sample distribution and $p_{1}(x)$ corresponds to the desired product distribution defined in Eq. 2. To draw a sample from the product distribution $p_{1}(x)$, we first initialize a sample from $p_{T}(x)$ and iteratively run MCMC on the sample to sample from each intermediate distribution before we finally draw a sample from $p_{1}(x)$. Since intermediate distributions are smoother and easier to sample from, this enables us to more effectively find a high likelihood scene $x$ from the product distribution.
To construct intermediate probability distributions, in the continuous setting with diffusion (or flow) models, we can use the intermediate score functions learned by each composed diffusion model to define intermediate distributions. In particular, for each diffusion expert $p^{(i)}(x)$, the score functions $\{\nabla_{x}\log p_{t}^{(i)}(x)\}_{t=1}^{T}$ defines T separate distributions $p_{t}^{(i)}(x)$. We can then construct a sequence of annealed distributions $\{p_{t}(x)\}_{t=1}^{T}$, where
@@
and each distribution $p_{t}(x)$ is specified by the composed score function (Section 3.3), allowing us to use Langevin sampling to implement AIS of the product distribution.
While the main experiments on generative models in this paper use diffusion models, we can likewise define a sequence of intermediate distributions for other generative models. In the setting of autoregressive models, we can construct annealed distributions by considering the intermediate autoregressive marginals of the distribution $p(x_{1})$, $p(x_{1:2})$ and so forth, where $x_{1},x_{2}$ are data regions, or different tempered variants of the product distribution. We can then use Gibbs sampling to implement AIS of the product distribution.
@@
where each $x_{i}$ refers to the part of a scene $x$ represented by the generative expert $p^{(i)}(\cdot)$, and $x_{\text{pa}(i)}$ refers to the other parts of a scene specified by other experts the expert is conditioned on.
Making each generative expert conditionally dependent on the values of other experts reduces the multi-modality in the distribution $p^{(i)}(\cdot)$, enabling more efficient MCMC sampling on the product distribution. We empirically illustrate the benefits of such a conditional generative expert formulation in Fig. 2, where sampling from the product of conditionally dependent generative models enables more effective sampling from the final product distribution.
@@ 3.2 Parallel Sampling with Discriminative Experts
Algorithm 1 Product of Expert (PoE) Sampling (Sketch)
@@
5:         $x^{(l)}\leftarrow\mathcal{K}_{t\leftarrow t+1}(\cdot\mid x^{(l)})$triangleright MCMC initialization
@@
7:              $x^{(l)}\leftarrow\mathcal{K}_{t}(\cdot\mid x^{(l)})$triangleright MCMC step
@@
Next, we discuss how we can modify the annealed sampling procedure in Section 3.1 to sample from the full product distribution, including discriminative experts
$x\sim p(x):\propto\prod_{i=1}^{N}p^{(i)}(x)\prod_{j=1}^{M}q^{(j)}(x).$ (5)
Similar to the previous section, we can define a sequence of intermediate sampling distributions
@@
To implement AIS with these intermediate distributions, one simple approach is to use importance sampling. Similar to the previous section, we can use MCMC sampling to sample from the product of the intermediate generative expert distributions from Eq. 3. After generating a total of $L$ samples from this generative product distribution, we draw a weighted random sample from these $L$ samples, with each sample has an importance weight $\prod_{j=1}^{M}q^{(j)}_{t}(x)$, corresponding to the likelihood of the scene under the product of the discriminative experts.
An issue with this simple procedure is that each of the $L$ samples drawn from the MCMC may be correlated with each other, since MCMC is slow at mixing and may not cover the entire generative product distribution. To improve the coverage of the $L$ samples drawn from MCMC, we use a parallel SMC sampler doucet2001introduction , where we maintain $L$ particles over the course of annealed sampling. At each intermediate annealed distribution, we run MCMC on each particle to draw samples from the product of the generative distributions, and then weigh and resample particles based on their likelihood under the product of the discriminative experts.
Our overall algorithm for sampling from Eq. 5 requires only black-box access to each discriminator, requiring only knowledge of the likelihood a discriminator assigns to a sample $x$ and not additional information such as the gradient of the log-likelihood of the discriminator. For settings where we do have more detailed information about the form of a discriminative expert, we can then directly treat the discriminative expert as a generative expert, as in Section 3.1, and use approaches such as gradient-based MCMC sampling to efficiently sample from the composed expert distribution.
@@ 3.3 Framework Instantiation
Figure 1: Application on Image Object Insertion. Given a graphics engine configurations (top row), a textual description, and an input image (first column), our method produces visually plausible edits with objects inserted.
@@
Constructing Annealed Distribution with Flow Models.
We construct our annealed generative expert distributions $p_{t}(x)$ (Eq. 3) using flow models. Each $p_{t}^{(i)}(x)$ is generated with a predicted velocity $v^{(i)}_{t}(x)$, which can be converted to score $\nabla_{x}\log p_{t}^{(i)}(x)$ through linear transformations. To propagate a sample from one annealed distribution $p_{t+1}(x)$ to the next $p_{t}(x)$, we use one Euler step in the flow ODE using the composed velocity $v_{t}(x)=\sum_{i}v^{(i)}_{t}(x)$ of the generative expert distributions, corresponding to a transition kernel $\mathcal{K}_{t\leftarrow t+1}(x^{\prime}\mid x):=\delta(x^{\prime}-(x+v_{t}(x)\Delta t))$ where $\delta$ denotes the Dirac delta function and $\Delta t$ is the discretization step size. While this integration procedure does not draw a sample from $p_{t}(x)$, it serves as an effective initialization for proper MCMC sampling from $p_{t}(x)$. We then use Langevin dynamics on the composed score $s_{t}(x)=\sum_{i}\nabla_{x}\log p_{t}^{(i)}(x)$ to sample from $p_{t}(x)$, i.e., $\mathcal{K}_{t}(x^{\prime}\mid x):=\mathcal{N}(x^{\prime};x+\frac{\kappa^{2}}{2}s_{t}(x),\kappa^{2}I)$ with step size $\kappa$. Constructions with autoregressive models are deferred to Appendix C.
@@
We implement conditional generative experts $p_{t}^{(i)}(x_{i}\mid x_{\text{pa}(i)})$ by modifying the original generative experts $p_{t}^{(i)}(x_{i})$. In particular, we modify the generative flow $v^{(i)}_{t}(x_{i})$ to take into account the predicted flow at parent regions
@@
where $w$ is the learning rate. Justifications of this approximation are deferred to Section D.2.
@@
To define each annealed discriminative expert $q^{(j)}_{t}(x)$ with reward function $r^{(j)}(x)$, we define its intermediate counterpart $r_{t}^{(j)}(x):=r^{(j)}(\hat{x})$, where $\hat{x}$ is the endpoint of $x$ predicted using the flow model $v_{t}(x)$. We use pre-trained VLMs as reward functions in this work.
Figure 2: Baseline Comparisons on Image Object Insertion. Our method better adheres to input object pose configurations compared to RF-Solver. The 4-th column shows conditional sampling improves visual fidelity.
@@
RF-Solver $1.619$±$\pm 0.605$ $\underline{0.178}$±$\pm 0.061$ $0.518$±$\pm 0.168$ $\mathbf{0.618}$±$\pm 0.070$ $0.662$±$\pm 0.103$ $0.948$±$\pm 0.777$ $0.438$±$\pm 0.179$
Ours $\mathbf{1.429}$±$\pm 0.602$ $\mathbf{0.065}$±$\pm 0.019$ $\mathbf{0.827}$±$\pm 0.073$ $0.596$±$\pm 0.061$ $\mathbf{0.783}$±$\pm 0.082$ $\mathbf{1.175}$±$\pm 0.666$ $\mathbf{0.817}$±$\pm 0.129$
No Cond $\underline{1.511}$±$\pm 0.661$ $\mathbf{0.065}$±$\pm 0.018$ $\underline{0.727}$±$\pm 0.141$ $\underline{0.599}$±$\pm 0.059$ $\underline{0.718}$±$\pm 0.120$ $\underline{1.142}$±$\pm 0.668$ $\underline{0.675}$±$\pm 0.180$
Table 1: Image Object Insertion. Our method achieves better foreground, background, and textual alignment and comparable visual quality compared to RF-Solver; $\pm$ denotes standard deviation.

### v2

@@ Abstract
Modern neural models capture rich priors and have complementary knowledge over shared data domains, e.g., images and videos. Integrating diverse knowledge from multiple sources—including visual generative models, visual language models, and sources with human-crafted knowledge such as graphics engines and physics simulators remains under-explored. We propose a probabilistic framework that combines information from these heterogeneous models, where expert models jointly shape a product distribution over outputs. To sample from this product distribution for controllable image/video synthesis tasks, we introduce an annealed MCMC sampler in combination with SMC-style resampling to enable efficient inference-time model composition. Our framework empirically yields better controllability than monolithic methods and additionally provides flexible user interfaces for specifying visual generation goals.  [^1] Project page: https://product-of-experts.github.io/.
@@ 1 Introduction
Recent image and video generative models (Saharia et al., 2022; Rombach et al., 2022; Ho et al., 2022; Brooks et al., 2024) have achieved remarkable success in realistic appearance modeling, yet still have limitations in following complex text instructions and adhering to real-world constraints such as physical laws. The former would benefit from semantic priors in visual language models (VLMs) (Radford et al., 2021; Li et al., 2023; Bai et al., 2023; Achiam et al., 2023), and the latter from rules embedded in physics simulators. However, training a single model to absorb all information sources (texts/visual corpora, simulation trajectories, etc.) can be prohibitively expensive.
We address this challenge by integrating knowledge across models at inference time (Du & Kaelbling, 2024), leveraging visual generative priors, discriminative rewards from VLMs, and rule-based knowledge from physics simulators. We focus on the task of controllable visual generation and aggregate opinions across a set of “expert” models by sampling from the product distribution defined by the models. Each expert, in our case, pretrained models such as VLMs or image and video generative models, focuses on one or more constraints in generation, and the resulting product distribution naturally assigns high probabilities only to samples that satisfy all constraints simultaneously. The use of product distributions to combine the opinions of multiple experts has been extensively used in the past (Bacharach, 1972; Genest & Zidek, 1986; Hinton, 1999), and in the visual domain in Markov Random Forests (Wang et al., 2013; Kolmogorov & Zabih, 2002; Glocker et al., 2008; Boykov & Funka-Lea, 2006) and Conditional Random Fields (Boykov & Jolly, 2001; Kumar & Hebert, 2003; He et al., 2004).
Despite its conceptual appeal, sampling from product distributions is often intractable. Straightforward approaches like rejection sampling are inefficient due to vanishing acceptance rates in high dimensions. Recent works have gotten around these difficulties by either composing model distributions in a simpler Gaussian latent space (Huang et al., 2022) or by composing through models through an annealed Langevin procedure (Du et al., 2020; 2023; Geffner et al., 2023), where the combination of annealing and gradient-based sampling enables the modes of the product probability distribution to be effectively identified. In this paper, we provide a more general framework that enables us to sample from the product of both autoregressive and diffusion-based generative models, as well as discriminative models such as VLMs. Our approach relies on a combination of Annealed Importance Sampling (Neal, 2001) and Sequential Monte Carlo (Doucet et al., 2001a) to effectively find the modes of the product distribution across all experts.
Our overall framework enables us to integrate multiple input constraints without costly retraining. We empirically validate its benefits in image generation following complex text instructions. It also naturally yields a flexible user interface for defining complex generative goals by selecting and configuring experts, e.g., allowing users to specify the pose and motion trajectory of an object of interest, insert it into an existing image, and animate the full scene.
Overall, our contributions are threefold:
@@
    We formulate controllable visual synthesis tasks in a unified Product-of-Experts (PoE) framework that enables principled knowledge integration of heterogeneous generative and discriminative models, as well as physics simulators.
@@
    We propose a practical, computationally efficient sampling framework based on Annealed Importance Sampling (AIS) and Sequential Monte Carlo (SMC).
@@ 2 Related Works
Prior works on compositional generative modeling typically combine multiple generative models to jointly generate data samples (Du et al., 2020; Garipov et al., 2023; Du et al., 2023; Huang et al., 2022; Mahajan et al., 2024; Du & Kaelbling, 2024; Bradley et al., 2025; Thornton et al., 2025; Gaudi et al., 2025) and has been previously applied in visual content generation (Liu et al., 2022a; Bar-Tal et al., 2023; Zhang et al., 2023; Yang et al., 2024; Su et al., 2024; Li et al., 2022), but such applications typically focus on multiple homogeneous generator experts, or, in the case of Li et al. (2022), a single generator expert with multiple discriminative experts. In contrast, our approach provides a general probabilistic framework through which many heterogeneous generative and discriminative experts can be jointly combined in generation.
@@
Our work is further related to recent work steering generative models with discriminative models and reward functions. Methods using gradient descent on a discriminative reward (Grathwohl et al., 2019; Dhariwal & Nichol, 2021; Bansal et al., 2023; Luo et al., 2025; He et al., 2023; Ye et al., 2024; Song et al., 2023; Rout et al., 2025) assume dense, differentiable gradients on the reward. Another branch of work applies SMC for more exact reward steering (Wu et al., 2023; Zhao et al., 2024; Singhal et al., 2025). Both Skreta et al. (2025); He et al. (2025) and our work build on top of this formulation and consider composing both generative and discriminative experts, but with the following difference: when considering generative expert products (Eq. 2), these works calculate path-wise importance weights to ensure sampling correctness, risking weight degeneracy as path length grows (Skreta et al., 2025), while ours uses per-timestep MCMC kernels that leave $p_{t}$ (Eq. 3) invariant without accumulating weights.
@@
Our framework integrates knowledge from physical simulators to improve physical accuracy compared to end-to-end models, while alleviating the need for tedious 3D scene setups in traditional graphics rendering pipelines. Recent works have explored such direction typically convert physical simulator outputs into a specific type of conditional signals such as optical flow (Burgert et al., 2025; Montanaro et al., 2024; Yang et al., 2025) and point tracking (Gu et al., 2025), or rely on a single pre-trained video generation model (Liu et al., 2024a; Chen et al., 2025; Tan et al., 2024), while our framework allows for the integration of various signals available from physical simulation, providing more complete specifications of controls.
3 Method
We aim to generate complex scenes $x\in\mathbb{R}^{d}$ by combining the priors from both generative models (such as image/video models conditioned on texts or control signals converted from physics simulators) as well as discriminative models (such as VLMs), with examples detailed in Section 4. We formulate this composition of models in a probabilistic manner, where generative models are represented as probabilistic priors over data $\{p^{(i)}(x)\}_{i=1}^{N}$. Discriminative models are written as constraint functions, $\{r^{(j)}:\mathbb{R}^{d}\rightarrow\mathbb{R}\}_{j=1}^{M}$, each assigning a scalar reward that represents how much a sample $x$ matches the constraint encoded in the model. Each constraint function is converted into an unnormalized probability distribution through the Boltzmann distribution, $q^{(j)}(x):=(\exp{{r^{(j)}}})(x)$.
To combine these two classes of models, we aim to sample from the product distribution:
@@ 3 Method
where each generative expert $p^{(i)}(x)$ and discriminative expert $q^{(j)}(x)$ are defined over parts of the scene $x$. The product distribution $p(x)$ has high probability for a sample $x$ when $x$ has high probability under both $p^{(i)}(x)$ and $q^{(j)}(x)$ (Hinton, 2002; Du et al., 2020; Du & Kaelbling, 2024), allowing us to compose together the priors in both generative and discriminative models.
However, sampling from Eq. 1 is challenging, as we need to search for samples that satisfies constraints across all experts. Below, we will discuss how we can efficiently and effectively sample from the Eq. 1 for high fidelity visual generation. We first discuss how we can sample from multiple generative experts in Section 3.1 and then discuss how to sample jointly from multiple discriminative and generative experts in Section 3.2, followed by a practical implementation in Section 3.3.
@@ 3.1 Sampling from Generative Experts
To sample from the above distribution, we can use Markov chain Monte Carlo (MCMC) Robert et al. (1999), where we iteratively refine samples based on their likelihood under the product distribution. In discrete settings, a variant of Gibbs sampling may be used, while in continuous domains, Langevin sampling based on the gradient of log density may be used (Welling & Teh, 2011).
A central issue, however, is that MCMC is a local refinement procedure and can take an exponentially long time for the sampling procedure to mix and find a high likelihood scene $x$ from the product distribution (Robert et al., 1999). To effectively sample from product of generative experts, we propose to use Annealed Importance Sampling (AIS) (Neal, 2001) and construct of a sequence of $T$ distributions, $\{p_{t}(x)\}_{t=T}^{1}$, where $p_{T}(x)$ is a smooth, easy-to-sample distribution and $p_{1}(x)$ is the desired product distribution defined in Eq. 2. To draw a sample from $p_{1}(x)$, we first initialize a sample from $p_{T}(x)$ and iteratively run MCMC on the sample to sample from each intermediate distribution before finally reaching $p_{1}(x)$. Since intermediate distributions are easier to sample from, this enables us to more effectively find a high likelihood scene $x$ from the product distribution.
We construct the intermediate probability distributions of the following form:
@@
where $p^{(i)}_{t}(x)$ is an expert-specific distribution interpolating between an initial distribution (e.g., Gaussian or uniform) and the final expert distribution $p^{(i)}(x)$. We list concrete choices of $p^{(i)}_{t}(x)$ for two common classes of generative models in Section 3.3. For autoregressive models (Oord et al., 2016; Larochelle & Murray, 2011), it is the marginal distribution for a prefixed data region determined by $t$; for diffusion/flow models (Liu et al., 2022b; Lipman et al., 2022; Ho et al., 2020; Sohl-Dickstein et al., 2015), it is the time marginal distribution with estimated score functions.
@@
where each $x_{i}$ refers to the part of a scene $x$ represented by the generative expert $p^{(i)}$, and $x_{\text{pa}(i)}$ refers to the other parts of a scene specified by other experts the expert is conditioned on.
Making each generative expert conditionally dependent on the values of other experts reduces the multi-modality in the distribution $p^{(i)}(x)$, enabling more efficient MCMC sampling on the product distribution. Its benefit is empirically shown in Fig. 2 and Table 1.
@@ 3.2 Parallel Sampling with Discriminative Experts
Algorithm 1 Product Distribution Sampling
@@
5:   $x^{(l)}\leftarrow\mathcal{K}_{t\leftarrow t+1}(\cdot\mid x^{(l)})$ $\triangleright$ MCMC initialization
@@
7:     $x^{(l)}\leftarrow\mathcal{K}_{t}(\cdot\mid x^{(l)})$ $\triangleright$ MCMC step
@@
Next, we discuss how we can modify the annealed sampling procedure in Section 3.1 to sample from the full product distribution from Eq. 1, including discriminative experts. Similar to the previous section, we can define a sequence of intermediate sampling distributions
@@
To implement AIS, one simple approach is importance sampling as follows. We first obtain $L$ samples from the product of the intermediate generative expert distributions defined in Eq. 3 according to Section 3.1, and then draw a weighted random sample from these $L$ samples, each with importance weight $\prod_{j=1}^{M}q^{(j)}_{t}(x)$ corresponding to the likelihood under product of discriminative experts.
An issue with this simple procedure is that each sample drawn from the MCMC may be correlated with each other, since MCMC is slow at mixing and may not cover the entire generative product distribution. To improve coverage, we use a parallel SMC sampler (Doucet et al., 2001b), where we maintain $L$ particles over the course of annealing. At each intermediate distribution, we run MCMC on each particle to draw samples from the product of the generative distributions, and then weigh and resample particles based on their likelihood under the product of the discriminative experts.
Our overall algorithm requires only black-box access to each discriminator, requiring knowing only the likelihood the discriminator assigns to a sample $x$ and not additional information such as log-likelihood gradients. If we do have more information about the discriminative expert’s form, we can directly treat it as generative as in Section 3.1 and use approaches such as gradient-based MCMC.
Figure 1: Application on Image Object Insertion where the goal is to insert assets posed in a graphics engine (top row) and described with text prompts (bottom row) into images (first column).
@@ 3.3 Framework Instantiation
Generative Experts.
The framework requires specifying the annealed generative distributions factors $\{p_{t}^{(i)}(x)\}_{t=T}^{1}$ (Eq. 3), a kernel $\mathcal{K}_{t\leftarrow t+1}(\cdot\mid\cdot)$ to propagate samples from the previous annealed distribution $p_{t+1}(x)$ to the next distribution $p_{t}(x)$ (line 5, Algorithm 1), and a MCMC kernel $\mathcal{K}_{t}(\cdot\mid\cdot)$ that leaves $p_{t}$ invariant (line 7). While $\mathcal{K}_{t\leftarrow t+1}$ does not draw a sample from $p_{t}(x)$ exactly, it serves as an effective initialization for MCMC sampling.
When generative expert $p^{(i)}(x)$ is a flow model, a natural choice for $\{p_{t}^{(i)}(x)\}_{t=T}^{1}$ is a discretization of the probability path generated by its velocity prediction $v^{(i)}_{t}(x)$. We set $\mathcal{K}_{t\leftarrow t+1}$ as one Euler step in the flow ODE integration, and $\mathcal{K}_{t}$ is the Langevin dynamics transition, both under composed score $\sum_{i}\nabla_{x}\log p_{t}^{(i)}(x)$ (Section C.1). When $p^{(i)}$ is an autoregressive model, we use its prefix marginal for annealing, namely $p_{t}^{(i)}(x)=p^{(i)}(x_{1:T+1-t})$ such that $x_{1:T}=x$, where $x_{1:T+1-t}$ is a data slice. We set $\mathcal{K}_{t\leftarrow t+1}$ such that it appends the next data slice sampled from the model-predicted conditional distribution, and $\mathcal{K}_{t}$ can be implemented as Gibbs sampling (Section C.2).
@@
We implement conditional generative experts $p_{t}^{(i)}(x_{i}\mid x_{\text{pa}(i)})$ for flow models by modifying the original generative experts $p_{t}^{(i)}(x_{i})$. In particular, we modify the generative flow $v^{(i)}_{t}(x_{i})$ to take into account the predicted flow at parent regions
@@
where $w$ is the learning rate. Justifications of this approximation are deferred to Section C.3.
@@
[…]
Figure 2: Image Object Insertion Comparisons. Our method better adheres to input geometric conditions while faithfully preserving background details. The last column shows that conditional sampling improves visual harmonization and fidelity.
@@
Graphics Engine Rendering Input
[…]
Ours No Cond $\underline{1.511}$$\pm 0.661$ $\mathbf{0.065}$$\pm 0.018$ $\underline{0.727}$$\pm 0.141$ $\underline{0.599}$$\pm 0.059$ $\underline{0.718}$$\pm 0.120$ $\underline{1.142}$$\pm 0.668$ $\underline{0.675}$$\pm 0.180$
[…]
Character Rendering Input (Magic Insert (Ruiz et al., 2024) Dataset)
[…]
Table 1: Image Object Insertion Evaluation.
Figure 3: Application on Physical-Simulation-Instructed Video Generation. Given an input image and a physical simulator rendering describing precise object motions, our method generates videos aligned with input motions while synthesizing natural content for non-foreground regions.